A company wants to reach a service level of 90%, which Z=1.28. The mean demand during lead time is 50 units with a standard deviation of 5. The ROP would be
Answer #1

Z=(X-µ)/σ

Here Z = 1.28

µ = 50

σ = 5

X = (1.28 * 5 )+ 50 = 6.4 + 50 = 56.4  

ROP = 57 units(Rounding to zero for units)

Please revert back in case of any doubt.

Please upvote. Thanks in advance.